# A Review on Single-Problem Multi-Attempt Heuristic Optimization

### v1

@@ Abstract.
In certain real-world optimization scenarios, practitioners are not interested in solving multiple problems but rather in finding the best solution to a single, specific problem. When the computational budget is large relative to the cost of evaluating a candidate solution, multiple heuristic alternatives can be tried to solve the same given problem, each possibly with a different algorithm, parameter configuration, initialization, or stopping criterion. The sequential selection of which alternative to try next is crucial for efficiently identifying the one that provides the best possible solution across multiple attempts. Despite the relevance of this problem in practice, it has not yet been the exclusive focus of any existing review. Several sequential alternative selection strategies have been proposed in different research topics, but they have not been comprehensively and systematically unified under a common perspective.
This work presents a focused review of single-problem multi-attempt heuristic optimization. It brings together suitable strategies to this problem that have been studied separately through algorithm selection, parameter tuning, multi-start and resource allocation. These strategies are explained using a unified terminology within a common framework, which supports the development of a taxonomy for systematically organizing and classifying them.
@@
Preprint submitted to ACT TELO July 22, 2025.
journal: TELOjournalvolume: Preprint submitted, July 22, 2025ccs: Mathematics of computing Mathematical optimizationccs: Computing methodologies Heuristic function constructionccs: Information systems Decision support systemsccs: Human-centered computing Interaction design process and methodsccs: General and reference Surveys and overviews
@@ 1. Introduction
An optimization problem consists of finding the best solution from a search space that either maximizes or minimizes a given objective function, subject to specific constraints (Arora, 2015). These problems are encountered in a wide range of disciplines, including logistics, scientific research, and engineering. In some cases, the objective is to solve multiple related problems, each defined by variations in the searh space and the objective function. Gago-Carro et al. (2024) address a location-assignment problem to find the ambulance service management that maximizes the number of emergency calls attended under different demand scenarios. In other situations, the goal is to solve a single, well-defined problem. This setting is common in real-world applications where the interest lies in optimizing a specific search space and objective function. For example, in computational biology, Grado et al. (2018) aim to identify the set of brain stimulation parameters that minimizes beta power in patients with Parkinson’s disease. In renewable energy systems, Reddy (2020) searches for the wind farm layout that maximizes annual energy production based on a given terrain and wind conditions, while Zarketa-Astigarraga et al. (2023) optimize the design of a Wells turbine to maximize its efficiency across a selected set of representative sea states. In this work, we focus on single-problem scenarios.
@@
There are different strategies in the literature that address the selection of the heuristic method to be considered in each attempt to solve a single optimization problem. Some works focus on restart strategies for a single local search algorithm (López-Sánchez et al., 2014; Mathesen et al., 2021), which allows solving the same problem in multiple attempts, considering different initializations of the algorithm at each one. Luersen and Le Riche (2004) focus on the kernel density estimation-based restart of the Nelder-Mead algorithm. Auger and Hansen (2005) instead, implement a restart strategy for the Covariance Matrix Adaptation Evolution Strategy, where the initial population is resampled uniformly in the search space at each restart, and the population size is progressively increased following a predefined schedule. This allows executing each attempt not only with a newly sampled initial population, but also with a systematically varied population size parameter, which is known to significantly influence the behavior and performance of evolutionary algorithms. Other works consider the joint selection of multiple algorithms along with their configurations. Souravlias et al. (2019) address a combinatorial optimization problem by defining a strategy to distribute the available computational budget across multiple attempts, each corresponding to a different configuration of Tabu Search, Variable Neighbourhood Search, or Multistart Local Search algorithms. Finally, we identify successive halving-based strategies (Li et al., 2018; Das Dores et al., 2018), which primarily focus on defining stopping criteria for each attempt. These approaches aim to allocate more evaluations to parameter configurations of supervised learning algorithms that show greater potential, in order to converge toward the best-performing configuration that minimizes the empirical error.
@@
The above examples show a wide diversity of strategies for addressing a single problem in multiple attempts. The diversity lies not only in the selection criteria, but also in the component that differentiates the heuristic methods selected for each attempt: a heuristic algorithm, a configuration of its parameters, an initialization or a stopping criterion. Thus, each combination of these four components defines a heuristic alternative to solve the optimization problem of interest. We have observed that the alternative selection strategies focus on solving a single problem are divided into four different topics studied separately in the literature, each one focused mainly on the selection of one component. However, all these strategies are valid for the same practical Single-Problem Multi-attempt Heuristic Optimization (SIMHO) scenario, where a practitioner faces the following question: how to sequentially select heuristic alternatives to find the best possible solution to a given optimization problem in multiple attempts?. There are reviews that give an overview of alternative selection (Kotthoff, 2016; Martí et al., 2019). However, most works focus on the selection of a single component (Locatelli and Schoen, 2016; Huang et al., 2020a), while only a few address the joint selection of two components at most (Kerschke et al., 2019; Yang and Shami, 2020; Bischl et al., 2023).
@@
Existing reviews do not comprehensively cover strategies for practitioners facing specific SIMHO problems. To overcome this gap, this paper presents a review focused on selection strategies suitable for SIMHO, encompassing strategies that address the selection of any combination of the four components. We unify existing literature related to SIMHO within a single framework, identifying a shared structure among strategies typically studied separately and formalizing them under the same terminology. Furthermore, we define a taxonomy that characterizes all strategies within the proposed framework. To the best of our knowledge, this is the first review to unify, describe, and classify selection strategies exclusively for SIMHO while considering all four components together.
@@
The rest of the paper is organized as follows. Section 2 classifies related alternative selection strategies, identifies in that classification those valid for SIMHO, and highlights the difference of the review proposed in this work from the perspective that already exists. Section 3 introduces the framework that sets the notation, formalizes SIMHO as an alternative selection problem, and defines the abstraction that generalizes within the same terminology the existing suitable strategies. Section 4 describes the four topics in the literature about alternative selection that address SIMHO, illustrating how all of them fit into the proposed abstraction. Section 5 defines the taxonomy and characterizes the different strategies as particular cases of the framework. Section 6 discusses the utility of the proposed framework-based review. Finally, Section 7 concludes the paper.
@@ 2. Related work
A major limitation of multi-problem strategies is that they must invest a significant number of attempts to build accurate models by solving multiple related instances prior to addressing the target problem, which often results in a time-consuming process (Eiben et al., 2007). Furthermore, their effectiveness depends on access to a sufficiently large and representative set of instances (Baudiš and Pošík, 2014; Kerschke et al., 2019), which is crucial for ensuring generalization to unseen problems. This may not be feasible in settings such as SIMHO, where only a specific optimization instance is available or the focus. Nevertheless, the multi-problem strategies remain highly valuable in contexts where users must routinely address problems belonging to the same family (Roman et al., 2016).
@@
Overall, each suitable strategy for SIMHO may focus on the selection of different components. However, existing reviews in the literature typically cover at most the selection of two out of the four components (see Table 1), as their primary aim is not to unify independently studied strategies across the different topics relevant to SIMHO. Instead, they provide broader overviews of alternative selection strategies that often include related but misaligned approaches, such as multi-problem strategies that do not apply to the SIMHO setting. For instance, Huang et al. (2020a) revise single-problem strategies for parameter configuration but also include offline parameter tuning methods, which fall outside the SIMHO scope. As a result, such reviews often lack a unified formalization and rely on fragmented taxonomies with component-specific descriptions. In contrast, this review focuses exclusively on single-problem multi-attempt strategies and introduces a common framework that formalizes the strategies using consistent terminology. This unified perspective enables a more coherent and comprehensive taxonomy, providing a clearer and more practical reference for researchers and practitioners working specifically on SIMHO, while avoiding the inclusion of strategies that are not directly applicable.
Table 1. Unification of the literature to address SIMHO.
Reviews Components
Heuristic algorithm Parameter configuration Initialization Stopping criterion
(Kotthoff, 2016) (Souravlias et al., 2021) (Kerschke et al., 2019)
@@
(Martí et al., 2019) (Locatelli and Schoen, 2016)
(Yang and Shami, 2020) (Bischl et al., 2023) (Elshawi et al., 2019)
@@ 3. SIMHO framework
An optimization problem consists of finding the best solution among a set of possible ones, which maximizes a specific performance function. Formally, it could be defined as follows
@@
where $X$ is the solution space and $f$ is the objective function that measures how good a solution is [^2] Without loss of generality, we assume that the optimization problem is a maximization problem. In the case of minimization, it would be enough to replace $f(x)$ in Equation (1) with f(x)-f(x).. In this section, we define the conditions considered to solve Problem (1), and the generalization of the suitable literature strategies to address it.
@@ 3.1. Multi-attempt heuristic optimization
where the alternative of the highest possible quality must be found over the multiple available executions. Thus, the reformulation shifts the exploration of the solution space $X$ to the set of alternatives $A$. The information of the same $A$ varies on different optimization problems, thus the qualities of each alternative $a\in A$ will be different for different definitions of Problem (1). Therefore, finding the best $a\in A$ to solve Problem (4) is in itself an optimization problem, where the desired solution is the best-visited solution $x_{a}$ provided by the $a\in A$ with the highest quality $f_{a}$. We define this problem generically and any approach that addresses it as SIMHO problem and strategy, respectively.
@@ 3.2. The abstract SIMHO strategy
where $p$, $p_{H}$, $p_{\Theta}$, $p_{\Xi}$ and $p_{C}$ denote the probability distributions over the set of alternatives $A$ and set of components $H$, $\Theta$, $\Xi$ and $C$, which we call distribution of alternatives and component distributions, respectively. Notice that to define the probability of $a$, its factors could use the informations from $\mathcal{I}$ associated with the executed alternatives with the same heuristic algorithm, or the same algorithm and parameter configuration, or the same algorithm, configuration and initialization (see Appendix A for a more detailed description).
@@
1 $\mathcal{I}\leftarrow\emptyset$;
@@ 7  $p\leftarrow$ Update probability distribution over $A$ using $\mathcal{I}$;
10return $\widetilde{x}=\operatorname*{\arg\max}_{x_{a}\in\mathcal{I}}f_{a}$
@@
We propose an abstract definition that formalizes the existing SIMHO strategies proposed in separately studied topics under the same terminology. This consists of an iterative alternation of two processes: the information generation and the distribution update. The first focuses on expanding the observed information $\mathcal{I}$ to enable the second to update the distribution of alternatives $p$ based on it. The pseudocode associated with our proposal is shown in Algorithm 1. It starts by initializing $p$ (line 2). Then, until the available budget is exhausted, in each iteration the alternation of the two processes is applied. Firstly, to generate new information, an unexecuted alternative $a$ is sampled from $A$ based on its current distribution (line 4). Then, the selected alternative is executed to reveal its information $I_{a}$ (line 5) and add this new information to the observed information $\mathcal{I}$ (line 6). Secondly, the distribution is updated by reassigning to each unexecuted alternative a probability based on the observed information (line 7). Finally, when no more time is left to carry out a new iteration, the best-visited solution with the highest objective value among all the executed alternatives is returned (line 9) [^4] Abusing the notation, we write $x_{a}\in\mathcal{I}$ to refer solely to the third elements (best-visited solutions) of the triplets (3) that form the observed information $\mathcal{I}$. In Sections 5.3.1 and 5.3.2, the same notation is used to refer solely to the observed execution times ($t_{a}\in\mathcal{I}$) and the observed qualities ($f_{a}\in\mathcal{I}$). .
@@ 4. Topics addressing SIMHO
We have observed that the strategies suitable to address SIMHO are proposed within four different topics related to alternative selection. All of them address optimization problems in multiple attempts with a set of alternatives, and the strategies proposed in them focus on determining the alternative to be executed in each attempt until the available executions are exhausted. Specifically, each topic names the iterative process [^5] The same or similar terms are used in the literature to refer to different topics that address different objectives, as as has been shown throughout Section 2. It should be emphasized, that we chose the terms that best represent the strategies addressed by each topic, although they may be terms used for other objectives in the literature as well. depending on which of the four components of the alternative it determines. To illustrate how the proposed abstraction generalizes the existing strategies in the four topics, we describe each of them both in general terms and using illustrative examples within Algorithm 1. For this purpose, we identify in them: an optimization problem, a set of alternatives, multiple attempts and the two alternating processes.

### v2

@@ Abstract.
In certain real-world optimization scenarios, practitioners are not interested in solving multiple problems but rather in finding the best solution to a single, specific problem. When the computational budget is large relative to the cost of evaluating a candidate solution, multiple heuristic alternatives can be tried to solve the same given problem, each possibly with a different algorithm, parameter configuration, initialization, or stopping criterion. In this practically relevant setting, the sequential selection of which alternative to try next is crucial for efficiently identifying the best possible solution across multiple attempts. However, suitable sequential alternative selection strategies have traditionally been studied separately across different research topics and have not been the exclusive focus of any existing review. As a result, the state-of-the-art remains fragmented for practitioners interested in this setting, with surveys either covering only subsets of relevant strategies or including approaches that rely on assumptions that are not feasible for the single-problem case.
This work addresses the identified gap by providing a focused review of single-problem multi-attempt heuristic optimization. It brings together suitable strategies for this setting that have been studied separately through algorithm selection, parameter tuning, multi-start, and resource allocation. These strategies are described using a unified terminology within a common framework, which supports the construction of a taxonomy for systematically organizing and classifying them. The resulting comprehensive review facilitates both the identification and the development of strategies for the single-problem multi-attempt setting in practice.
@@
© [Judith Echevarrieta, Etor Arza, Aritz Pérez, Josu Ceberio] [2026]. This is the author’s version of The Work. It is posted here for your personal use. Not for redistribution. The definitive version was published in ACM Transactions on Evolutionary Learning and Optimization, https://doi.org/10.1145/3795878.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXjournal: TELOccs: Mathematics of computing Mathematical optimizationccs: Computing methodologies Heuristic function constructionccs: General and reference Surveys and overviews
@@ 1. Introduction
An optimization problem consists of finding the best solution from a search space that either maximizes or minimizes a given objective function, subject to specific constraints (Arora, 2015). These problems are encountered in a wide range of disciplines, including logistics, scientific research, and engineering. In some cases, the objective is to solve multiple related problems, each defined by variations in the search space and the objective function. Gago-Carro et al. (2024) address a location-assignment problem to find the ambulance service management that maximizes the number of emergency calls attended under different demand scenarios. In other situations, the goal is to solve a single, well-defined problem. This setting is common in real-world applications where the interest lies in optimizing a specific search space and objective function. For example, in computational biology, Grado et al. (2018) aim to identify the set of brain stimulation parameters that minimizes beta power in patients with Parkinson’s disease. In renewable energy systems, Reddy (2020) searches for the wind farm layout that maximizes annual energy production based on a given terrain and wind conditions, while Zarketa-Astigarraga et al. (2023) optimize the design of a Wells turbine to maximize its efficiency across a selected set of representative sea states. In this work, we focus on the second single-problem scenario.
@@
There are different strategies in the literature that address the selection of the heuristic method to be considered in each attempt to solve a single optimization problem. Some works focus on restart strategies for a single local search algorithm (López-Sánchez et al., 2014; Mathesen et al., 2021), which allow solving the same problem in multiple attempts, considering different initializations of the algorithm at each one. Luersen and Le Riche (2004) focus on the kernel density estimation-based restart of the Nelder-Mead algorithm. Auger and Hansen (2005) instead, implement a restart strategy for the Covariance Matrix Adaptation Evolution Strategy, where the initial population is resampled uniformly in the search space at each restart, and the population size is progressively increased following a predefined schedule. This allows executing each attempt not only with a newly sampled initial population, but also with a systematically varied population size parameter, which is known to significantly influence the behavior and performance of evolutionary algorithms. Other works consider the joint selection of multiple algorithms along with their configurations. Souravlias et al. (2019) address a combinatorial optimization problem by defining a strategy to distribute the available computational budget across multiple attempts, each corresponding to a different configuration of Tabu Search, Variable Neighbourhood Search, or Multistart Local Search algorithms. Finally, we identify successive halving-based strategies (Li et al., 2018; Das Dores et al., 2018), which primarily focus on defining stopping criteria for each attempt. These approaches aim to allocate more evaluations to parameter configurations of supervised learning algorithms that show greater potential, in order to converge toward the best-performing configuration that minimizes the empirical error.
@@
The above examples show a wide diversity of strategies for addressing a single problem in multiple attempts. The diversity lies not only in the selection criteria, but also in the component that differentiates the heuristic methods selected for each attempt: a heuristic algorithm, a configuration of its parameters, an initialization or a stopping criterion. Thus, each combination of these four components defines a heuristic alternative to solve the optimization problem of interest. We have observed that alternative selection strategies focus on solving a single problem are divided into four different topics studied separately in the literature, each one focused mainly on the selection of one component. However, all these strategies are valid for the same practical Single-Problem Multi-attempt Heuristic Optimization (SIMHO) scenario, where a practitioner faces the following question: how to sequentially select heuristic alternatives to find the best possible solution to a given optimization problem in multiple attempts?. There are reviews that give an overview of alternative selection strategies (Kotthoff, 2016; Martí et al., 2019), but most either focus on limited subsets of relevant strategies for SIMHO (Locatelli and Schoen, 2016; Martí et al., 2019; Yang and Shami, 2020) or include approaches that rely on assumptions that make them not feasible or not efficiently applicable to the SIMHO setting (Huang et al., 2020a; Kerschke et al., 2019; Bischl et al., 2023).
@@
Existing reviews do not comprehensively cover strategies for practitioners facing specific SIMHO settings. To overcome this gap, this paper presents a review focused on selection strategies suitable for SIMHO, encompassing strategies that address the selection of any combination of the four components. We unify existing literature related to SIMHO within a single framework, identifying a shared structure among strategies typically studied separately and formalizing them under the same terminology. Furthermore, we define a taxonomy that characterizes all strategies within the proposed framework. To the best of our knowledge, this is the first review to unify, describe, and classify selection strategies exclusively for SIMHO while considering all four components together.
@@
The rest of the paper is organized as follows. Section 2 classifies related alternative selection strategies, identifies in that classification those valid for SIMHO, and highlights the difference of the review proposed in this work from the perspective that already exists. Section 3 introduces the framework that sets the notation, formalizes SIMHO as an alternative selection problem, and defines the abstraction that generalizes within the same terminology the existing suitable strategies. Section 4 describes the four topics in the literature about alternative selection that address SIMHO, illustrating how all of them fit into the proposed abstraction. Section 5 defines the taxonomy and characterizes the different strategies as particular cases of the framework. Section 6 discusses the utility and scope of the proposed framework-based review. Finally, Section 7 concludes the paper.
@@ 2. Related work
2.1. Single- and multi-problem strategies through the lens of SIMHO
@@
A major limitation of multi-problem strategies is that they must invest a significant number of attempts to build accurate models by solving multiple related instances prior to addressing the target problem, which often results in a time-consuming process (Eiben et al., 2007). Furthermore, their effectiveness depends on access to a sufficiently large and representative set of instances (Baudiš and Pošík, 2014; Kerschke et al., 2019), which is crucial for ensuring generalization to unseen problems. This may not be feasible in settings such as SIMHO, where only a specific optimization instance is available or the focus. Nevertheless, multi-problem strategies remain highly valuable in contexts where users must routinely address problems belonging to the same family (Roman et al., 2016).
@@
2.2. Existing reviews through the lens of SIMHO
Overall, each suitable strategy for SIMHO may focus on the selection of different components. There are multiple reviews in the literature that include single-problem selection strategies valid for SIMHO. However, they typically cover at most the selection of two or three out of the four components (see the third column of Table 1). This is because, in most cases, their primary aim is not to unify independently studied strategies across the different topics relevant to SIMHO. Instead, they provide broader overviews of alternative selection strategies that often include related but misaligned approaches, such as multi-problem strategies that do not apply to the SIMHO setting (see the second column of Table 1).
For instance, Souravlias et al. (2021) include algorithm portfolios and per-instance algorithm selection, Kotthoff (2016) differentiates between strategies based on dynamic and static features, Huang et al. (2020a) classify strategies as specialist or generalist, Schede et al. (2022) refer to instance-specific and offline configurators, Elshawi et al. (2019) describe both hyperparameter optimization and meta-learning strategies, or Bischl et al. (2023) which briefly discuss warm-start strategies for reducing computational time in hyperparameter optimization by transferring information from previous experiments on the problem of interest or similar problems to the one currently being considered. In all these cases, their dual focus fits our definitions of single-problem and multi-problem strategies, respectively. However, we compare our work with these reviews only through their parts dedicated to single-problem strategies, as the remaining parts fall outside the scope of SIMHO.
Table 1. Classification of existing reviews according to the type of selection strategies they focus on and the components addressed that are valid for SIMHO.
Reviews Focus on Components
Single-problem Multi-problem Heuristic algorithm Parameter configuration Initialization Stopping criterion
(Souravlias et al., 2021)
(Kotthoff, 2016)
@@ 2.2. Existing reviews through the lens of SIMHO
(Locatelli and Schoen, 2016) (Martí et al., 2019)
(Yang and Shami, 2020)
(Elshawi et al., 2019)
(Bischl et al., 2023)
@@
As a result of the dual focus of existing reviews, they often lack a unified formalization and rely on fragmented taxonomies with component-specific descriptions. In contrast, we provide a review that focuses exclusively on single-problem multi-attempt strategies, and we introduce a common framework that formalizes the strategies that address the selection of any combination of the four components using consistent terminology. This unified perspective enables a more coherent and comprehensive taxonomy, providing a clearer and more practical reference for researchers and practitioners working specifically on SIMHO, while avoiding the inclusion of strategies that are not efficiently applicable.
@@ 3. SIMHO framework
An optimization problem consists of finding the best solution among a set of possible ones, which maximizes a performance function. Formally, it could be defined as follows
@@
where $X$ is the solution space and $f$ is the objective function that measures how good a solution is [^2] Without loss of generality, we assume that the optimization problem is a maximization problem. In the case of minimization, it would be enough to replace $f(x)$ in Problem (1) with f(x)-f(x). In addition, as both the solution space $X$ and the objective function $f$ are specific, Problem (1) defines a single, well-defined optimization problem.. In this section, we define the conditions considered to solve Problem (1), and the generalization of the suitable literature strategies to address it.
@@ 3.1. Multi-attempt heuristic optimization
where the alternative of the highest possible quality must be found over the multiple available executions [^4] Despite the formal similarity to other existing problems (e.g., the classic algorithm selection introduced by (Rice, 1976) for multi-problem setting) this formulation is specific to single-problem settings because it depends on a single optimization problem determined by Problem (1). For a more detailed discussion, see Appendix B.. Thus, the reformulation shifts the exploration of the solution space $X$ to the set of alternatives $A$. From this point of view, finding the best $a\in A$ that solves Problem (4) becomes the new optimization problem, where the desired solution is the best-visited solution $x_{a}$ provided by the $a\in A$ with the highest quality $f_{a}$. We define this problem generically and any approach that addresses it as SIMHO problem and strategy, respectively.
@@ 3.2. The abstract SIMHO strategy
where $p$, $p_{H}$, $p_{\Theta}$, $p_{\Xi}$ and $p_{C}$ denote the probability distributions over the set of alternatives $A$ and the sets of components $H$, $\Theta$, $\Xi$ and $C$, which we call distribution of alternatives and component distributions, respectively. Notice that to define the probability of $a$, its factors could use the informations from $\mathcal{I}$ associated with the executed alternatives with the same heuristic algorithm, or the same algorithm and parameter configuration, or the same algorithm, configuration and initialization (see Appendix A for a more detailed description).
We propose an abstract definition that formalizes the existing SIMHO strategies proposed in separately studied topics under the same terminology. This consists of an iterative alternation of two processes: the information generation and the distribution update. The first focuses on expanding the observed information $\mathcal{I}$ to enable the second to update the distribution of alternatives $p$ based on it. The pseudocode associated with our proposal is shown in Algorithm 1. It starts by initializing $p$ (line 2). Then, until the available budget is exhausted, in each iteration the alternation of the two processes is applied. Firstly, to generate new information, an unexecuted alternative $a$ is sampled from $A$ based on its current distribution (line 4). Then, the selected alternative is executed to reveal its information $I_{a}$ (line 5) and add this new information to the observed information $\mathcal{I}$ (line 6). Secondly, the distribution is updated by reassigning to each unexecuted alternative a probability based on the observed information (line 7). Finally, when no more time is left to carry out a new iteration, the best-visited solution with the highest objective value among all the executed alternatives is returned (line 10) [^5] Abusing the notation, we write $x_{a}\in\mathcal{I}$ to refer solely to the third elements (best-visited solutions) of the triplets (3) that form the observed information $\mathcal{I}$. In Sections 5.3.1 and 5.3.2, the same notation is used to refer solely to the observed execution times ($t_{a}\in\mathcal{I}$) and the observed qualities ($f_{a}\in\mathcal{I}$). .
@@
1 $\mathcal{I}\leftarrow\emptyset$, $t\leftarrow 0$ ;
@@ 8  $t\leftarrow t+t_{a}$ ;
11return $\widetilde{x}=\operatorname*{\arg\max}_{x_{a}\in\mathcal{I}}f_{a}$
@@ 4. Topics addressing SIMHO
We have observed that the strategies that address SIMHO are proposed within four different topics related to alternative selection. All of them address specific optimization problems in multiple attempts with a set of alternatives, and the strategies proposed in them focus on determining the alternative to be executed in each attempt until the available executions are exhausted. Specifically, each topic names the iterative process [^6] The same or similar terms are used in the literature to refer to different topics that focus on different objectives, as has been shown throughout Section 2, and has been specified in the second column of Table 1. It should be emphasized that, although we tried to choose the terms that best represent the single-problem strategies addressed by each topic, these terms have other uses in the literature. depending on which of the four components of the alternative it determines. To illustrate how the proposed abstraction generalizes the existing strategies in the four topics, we describe each of them both in general terms and using illustrative examples within Algorithm 1. For this purpose, we identify in them: an optimization problem, a set of alternatives, multiple attempts and the two alternating processes.